A politician wants to estimate the proportion of citizens who "vote for him." What sample size should be obtained if he wants to be within 2 percentage points with 95% confidence if he uses an estimate of 14% obtained from a poll? Round up your answer. Use the z table from the canvas module 5.  
Answer #1

Solution :

Given that p = 0.14 , margin of error E = 0.02

=> q = 1 - p = 0.86

=> for 95% confidence level, Z = 1.96

=> Sample size n = p*q*(Z/E)^2

= 0.14*0.86*(1.96/0.02)^2

= 1156.3216

= 1156 (nearest integer)